### v2

# Continually Evolving Skill Knowledge in Vision Language Action Model

Abstract
Developing general robot intelligence in open environments requires continual skill learning. Recent Vision-Language-Action (VLA) models leverage massive pretraining data to support diverse manipulation tasks, but they still depend heavily on task-specific fine-tuning, revealing a lack of continual learning capability. Existing continual learning methods are also resource-intensive to scale to VLA models. We propose Stellar VLA, a knowledge-driven continual learning framework with two variants: T-Stellar, modeling task-centric knowledge space, and TS-Stellar, capturing hierarchical task–skill structure. Stellar VLA enables self-supervised knowledge evolution through joint learning of task latents and the knowledge space, reducing annotation needs. Knowledge-guided expert routing provide task specialization without extra network parameters, lowering training overhead. Experiments on the LIBERO benchmark and real-world tasks show over 50% average improvement in final success rates relative to baselines. TS-Stellar further excels in complex action inference, and in-depth analyses verify effective knowledge retention and discovery. Our code will be released soon.
1 Introduction
A core requirement for lifelong agents is continual learning that preserves prior knowledge while adapting to new situations. Vision-language-action (VLA) models have advanced robotic policy learning through large-scale pretraining [UniVLA, kim2024openvla]. However, unlike vision or language foundation models that generalize across tasks, they still rely heavily on task-specific fine-tuning [kim2025fine, li2025metavla, kawaharazuka2025vision], reflecting limited understanding of underlying skill relations. The diversity of manipulation skills and environmental complexity make continual learning a key challenge for VLA models toward general robotic intelligence.
Many prior works have explored continual imitation learning (CIL), where agents acquire new skills from demonstrations without forgetting old ones. However, most rely on task-specific adapters or handcrafted features [TAIL, SDP, wan2024lotus, IsCiL], limiting scalability to large VLA with many tasks and parameters. Recent task-centric representation methods [VQBET, mete2024quest, VQ-VLA] address this by improving cross-task generalization via generative modeling, yet their multi-stage training and manual annotation still hinder continual learning. LEGION [legion] further extends continual reinforcement learning with an infinite task knowledge space, offering insights for extending CIL to general VLA models.
Mixture-of-Experts (MoE) architectures [shazeer2017outrageously] have also shown strong scalability across multimodal tasks [switch, mixtral, deepseekmoe, wu2024deepseek, lin2024moe]. By dynamically routing among experts, MoE mitigates task interference and scales efficiently, potentially alleviating catastrophic forgetting [SDP]. These advances raise a key question: Can VLA uncover their latent task-relevant knowledge and leverage it for continual policy evolution through adaptive task specialization?
To explore this question, we propose Stellar VLA, an end-to-end framework for continual imitation learning in vision-language-based manipulation. Stellar VLA enables knowledge evolution through the mutual refinement of task-centric representations and a self-evolving knowledge space, forming a self-supervised cycle of knowledge retention and discovery. Inspired by LEGION [legion], we adopt a Dirichlet Process–based variational model to adaptively represent infinite task-relevant knowledge. By integrating task knowledge into expert routing in the action head, the model balances parameter sharing and specialization across tasks. We further introduce two variants: T-Stellar, focused on continual task-centric adaptation, and TS-Stellar, which additionally models hierarchical task–skill relations for reusable skill discovery.
We evaluate our approach in both simulation and real-world settings. On LIBERO benchmark, though trained from scratch, T-Stellar and TS-Stellar achieve strong overall performance, with average final success rates approximately 1.5× higher than those of the baselines [UniVLA, MoDE]. Real-world experiments further confirm these gains, where both variants outperform baselines and TS-Stellar, in particular, shows clear advantages in modeling complex hierarchical actions. Visualizations illustrate their ability to acquire new skills and preserve prior ones. To conclude our contribution:
  • We present Stellar VLA, a continual imitation learning framework that jointly learns task-centric representations and a self-evolving knowledge space. Two variants are introduced: T-Stellar for task-level modeling and TS-Stellar for hierarchical task–skill modeling.
  • We adopt a knowledge-guided expert routing mechanism, enabling balanced parameter sharing and specialization.
  • Experiments on LIBERO and real-world tasks show robust improvements in task adaptation, with both variants capturing evolving task or skill structure.
Figure 1: Overall architecture of Stellar VLA. CLIP [radford2021learning] and FiLM [perez2018film]-conditioned ResNet encode language and visual inputs respectively. The task-centric representation $z$ and knowledge space are jointly learned through knowledge update and latents aggregation, as detailed in Sec. $3.$ The learned knowledge prior finally guides the MoE action head for motion prediction, as detailed in Sec. 3.4.
2 Related Work
Continual Imitation Learning in Robotics. To address catastrophic forgetting in data-driven robotic policy learning, many studies explore Continual Imitation Learning (CIL) [IsCiL, Lifelong]. One line of research mitigates forgetting by introducing additional parameters for each new learning phase [rusu2016progressive, li2019learn, rao2019continual, mallya2018packnet, TAIL, SDP, IsCiL], while another relies on handcrafted task decomposition through clustering or multi-stage learning [IsCiL, wan2024lotus]. However, when scaled to large VLA models covering diverse tasks, the former causes parameter explosion, whereas the latter incurs heavy computation and preprocessing costs due to multimodal data complexity.
Task-centric Representation Learning.
Task-centric representation learning offers the potential for continual skill accumulation in robots. Prior studies [jiang2022efficient, jiang2023h, zheng2024prise, VQBET, mete2024quest] have explored action representation learning. VQ-VLA [VQ-VLA] recently extended this idea to large-scale VLA models, enabling unified action tokenization across modalities. However, these methods typically rely on two-stage training comprising action reconstruction and policy learning, resulting in high computational cost.
Recent studies [wang2024omnijarvis, wu2023unleashing, zhang2025dreamvla, hu2023thought, belkhale2024rt] also learn task-centric representations from visual observations and language goals. UniVLA [UniVLA] exemplifies this direction by predicting future frames and subgoals. Some work [VLA-OS, bridging] provide unified frameworks for analyzing representation strategies.  [bridging] proves vision-driven features excel in cluttered scenes, and VLA-OS [VLA-OS] takes an initial step toward continual learning with language and visual grounding. However, most methods still rely on task segmentation or grounding, requiring extensive manual annotation.
LEGION [legion] applies a Dirichlet Process Mixture Model [antoniak1974mixtures] for language task representation, forming an infinite knowledge space that enables strong continual reinforcement learning. Despite state priors, it highlights the potential of representation learning for cross-task retention.
In this work, we learn task-centric representations from vision-language data to unify perception and task intent. Motivated by LEGION, a self-evolving knowledge space is proposed to enable automatic knowledge acquisition, reducing dependence on manual annotations.
Policy Models with Mixture of Experts. Recent studies show that the Mixture-of-Experts (MoE) architecture [shazeer2017outrageously] effectively enhances model scalability and computation efficiency in both large language [switch, mixtral, deepseekmoe] and vision-language models [wu2024deepseek, lin2024moe, team2025kimi]. Recently, VLA models have also incorporated MoE structures to enhance task scalability. MoDE [MoDE] first introduces a diffusion-based MoE policy, where expert routing is determined by noise level rather than task semantics. Many subsequent studies [zhang2025mole, yu2025forcevla, zhao2025more, expertise, yang2025drivemoe] leverage MoE for task specialization and improved multitask performance, yet their exploration of continual learning remains limited. SDP [SDP], though not leveraging language information, demonstrates the potential of MoE structures for continual learning by introducing task-specific routers to reduce parameter interference. However, its experts mainly operate in perception modules and require additional expert layers for incremental adaptation. In this work, we propose a fixed-size task-specific MoE action head guided by priors from knowledge space, leveraging the strengths of MoE to enhance continual learning.
3 Method
In this section, we present the architecture of Stellar VLA (Fig. 1), which enables CIL for vision-language manipulation. Our goal is to continually evolve task-relevant knowledge, enabling Stellar VLA to acquire new skills while mitigating forgetting of prior tasks. We first define the CIL setting, then describe the design and evolution of the knowledge space, and finally explain the process of guiding the expert policy via knowledge priors.
3.1 Problem Formulation
In the CIL setting [IsCiL], an agent learns sequentially from a stream of tasks $\{\mathcal{T}_{j}\}_{j=1}^{\infty}$, where each task $j$ provides $N$ expert demonstrations $\{\tau_{j}^{i}\}_{i=1}^{N}$, consisting of language instructions $\boldsymbol{l}$, observations $\boldsymbol{o}$, and actions $\boldsymbol{a}$. Unlike multi-task learning assuming full access to all demonstrations, CIL requires acquiring new task knowledge under limited access to past data while retaining previously learned skills, reflecting the realistic demands of lifelong robot learning.
To alleviate forgetting caused by limited access to past tasks, following LOTUS [wan2024lotus], we adopt Experience Replay (ER) with a small buffer $\mathcal{B}$. After completing task $j$, $q\%$ of its demonstrations are stored in $\mathcal{B}$ and jointly trained with current data when learning task j+1j{+}1. However, relying on only a small portion of past data often leads to prediction drift on previous tasks $\mathcal{T}_{k}$ (k<j+1k<j{+}1). To address this, we design a VLA architecture that continually discovers and preserves task knowledge, enabling dynamic expert composition for better parameter sharing and reduced interference.
3.2 Dirichlet-Process-based Knowledge Space
To capture both shared and distinct patterns across tasks, merely learning task-centric representations [VQBET, mete2024quest, VQ-VLA] is insufficient. Latent representations describe task-relevant variations after filtering sensory noise, but remain low-level abstractions without explicit task affiliation, making it difficult to separate new from past tasks and often leading to catastrophic forgetting during continual learning. To obtain discrete and interpretable task- or skill-level knowledge, these representations are further organized into a high-level, dynamically evolving knowledge space.
In CIL, task-relevant knowledge would incrementally grow as new task data arrives. Finite parametric models, such as Gaussian Mixture Models [bing2022meta], cannot adaptively fit the expanding knowledge space. Following LEGION [legion], we adopt a Dirichlet Process (DP)–based nonparametric model that allows infinite mixture clustering.
Dirichlet-Process Model. In DP, the sampling process $G\sim DP(\alpha,G_{0})$ can be written as:
$G=\sum_{k=1}^{\infty}\pi_{k}\,\delta_{\theta_{k}^{*}},\quad\theta_{k}^{*}\sim G_{0},$ (1)
Here, $G$ is a probability measure and $G_{0}$ the base distribution. $\theta_{k}^{*}$ are distribution parameters, while $\pi_{k}$ are infinite-dimensional weights [sethuraman1994constructive]. Considering $G_{0}$ as the aggregated knowledge of infinite tasks, then $G$ could represent a random measure over the current task set. The discreteness of $G$ causes samples $\theta$ from $G$ to repeat, naturally forming clusters. Moreover, its nonparametric property allows new clusters to emerge as needed, making the knowledge space inherently open and self-evolving.
Building on DP-based clustering, we design two knowledge spaces: 1) a task-centric space using Dirichlet Process Mixture Model (DPMM) to form task clusters from the latent representations, and 2) a hierarchical task–skill space using a Hierarchical Dirichlet Process (HDP) to capture task–subskill relations.
Task-Centric Knowledge Modeling from DPMM. Task distributions in latent space are dynamically clustered using Dirichlet Process Mixture Model (DPMM), a Bayesian nonparametric model supporting potentially infinite clusters.
In DPMM, based on DP sampling in Eq. 1, the full generative process for $z_{j}$ in task $j$ under the task-set probability measure $G$ can be expressed as $z_{j}\sim F_{task}(\theta_{j}),\theta_{j}\sim G$, where $\theta_{j}$ denotes the distribution parameters for task $j$ sampled from $G$, and $F_{\text{task}}(\theta_{j})$ represents its corresponding observation model. Thus, sampling from DPMM corresponds to drawing a latent observation from a task distribution. Assuming $F_{task}$ is Gaussian, the DPMM models a knowledge space comprising an unbounded set of task-centric Gaussian distributions based on latent task representations.
Hierarchical Task-Skill Knowledge Modeling from HDP. The DPMM-based knowledge space assumes independent task distribution from $z_{j}$. However, robotic tasks often share subskills, and DPMM alone may fail to capture these relations in complex or long-horizon settings. To model such hierarchy, we extend it with a Hierarchical Dirichlet Process (HDP) [teh2006hierarchical].
Specifically, each task $j$ is modeled as a probability measure over its subset of skills $\{\text{skill}_{i}\}$. Thus, by applying a second-level DP sampling on the task-set measure $G$, we could obtain skill distribution parameters $\theta_{ji}$ and corresponding latent skill samples $z_{ji}$ in task $j$ as:
$z_{ji}\sim F_{skill}(\theta_{ji}),\theta_{ji}\sim G_{j},G_{j}\sim DP(\gamma,G)$ (2)
Assuming $F_{skill}$ as a Gaussian distribution, each skill $i$ could be parameterized by $\theta_{i}^{skill}=(\mu_{i}^{skill},\sigma_{i}^{skill})$. The distribution parameters of each task, $\theta_{j}^{task}=(\mu_{j}^{task},\sigma_{j}^{task})$, can be derived from its associated skill set $\{{skill}_{i}\}$:
$\begin{split}\mu_{j}^{task}&=\sum_{\text{skill }i\in\text{task }j}\pi_{ji}\,\mu_{i}^{skill},\\ {\sigma_{j}^{task}}^{2}&=\sum_{\text{skill }i\in\text{task }j}\pi_{ji}\,\big[{\sigma_{i}^{skill}}^{2}+(\mu_{i}^{skill}-\mu_{j}^{task})^{2}\big].\end{split}$ (3)
$\pi_{ji}$ is the local mixture weight of skill $i$ in task $j$. According to Eq. 1, task-level probability measure $G_{j}$ would share atoms $\theta_{k}$ with $G$, but has different mixture weights. This formulation naturally models shared skills across tasks, forming a hierarchical task–skill knowledge space.
3.3 Knowledge Space Learning via Self-Evolution
This section presents the co-evolution of the knowledge space with Stellar VLA during CIL, resulting in two variants: T-Stellar and TS-Stellar, based on task-centric and task–skill hierarchical knowledge spaces respectively.
Following Sec. 3.2, we consider the dependency chain from sensor observations through latent representations to the knowledge space. To finally get stable knowledge features, a hierarchical variational inference is performed: task-centric latent representations are inferred via VAE [kingma2013auto] from vision-language data for new task learning, and task knowledge is inferred from them through DP-base models to manage old and new task information.
Task-Centric Representation Learning. To obtain semantically rich task-centric representations $z_{j}$ for task $j$, we employ a VAE that reconstructs vision–language inputs. Unlike standard VAE that learn latents from a predefined Gaussian prior, $z_{j}$ is constrained to aggregate within the task-relevant knowledge space as shown in Fig. 1, thus reducing forgetting of past experiences. Specifically, the model is trained by minimizing the reconstruction loss $L_{recon}$ and KL divergence $L_{kl}$, with task-specific adaptations for knowledge-space alignment.
In T-Stellar, the language goal $\boldsymbol{l}$ and observation $\boldsymbol{o}$ serves as the reconstruction target. $L_{kl}$ is computed from the probability $p_{jk}$ of $z_{j}^{task}$ belonging to Gaussian cluster $k$.
$L_{\text{KL}_{j}}=\sum_{k=1}^{K_{c}}p_{jk}L_{\text{KL}_{jk}}$ (4)
Here, $K_{c}$ is the number of clusters in the current DPMM, and $L_{\text{KL}_{jk}}$ is the KL divergence between $z_{j}$ and cluster $k$, measuring the similarity between the task-centric representation and DPMM components.
For TS-Stellar which hierarchically models tasks and skills, we use a hierarchical VAE to approximate their latent vectors. The task representation $z_{j}^{task}$ from the encoder is passed through linear layers along with visual embeddings to generate skill representation $z_{j}^{skill}$. The reconstruction loss $L_{recon}^{\text{TS}}$ is then given by:
$L_{\text{recon}}^{\text{TS}}=L_{\text{recon}}\!\big(f_{\text{dec}}^{\text{task}}(z_{j}^{\text{task}}),\boldsymbol{l}_{j}\big)+L_{\text{recon}}\!\big(f_{\text{dec}}^{\text{skill}}(z_{j}^{\text{skill}}),\boldsymbol{o}_{j}\big)$ (5)
which measures encoding of language $\boldsymbol{l}$ via $z_{j}^{task}$ and observation $\boldsymbol{o}$ via $z_{j}^{skill}$. Visual observations implicitly capture task-related skills. For $L_{kl}$, HDP defines task- and skill-level losses: task parameters are obtained from known skill clusters (Eq. 3), and $L_{kl}^{HDP}$ sums the KL divergences between $(z_{j}^{task},\{\theta^{task}\})$, and $(z_{j}^{skill},\{\theta^{skill}\})$:
$\mathcal{L}_{\text{KL}_{j}}^{\text{TS}}=\mathcal{L}_{\text{KL}_{j}}(z_{j}^{task},\theta^{task})+\mathcal{L}_{\text{KL}_{j}}(z_{j}^{skill},\theta^{skill})$ (6)
Knowledge Space Distribution Learning. As task-centric representations are continuously learned from vision–language observations under the constraints of DP models, the DP-based knowledge space is simultaneously updated based on these representations, ensuring timely discovery of new knowledge, as illustrated in Fig. 1. Considering $z_{j}$} as observations of the knowledge space, we perform variational inference to approximate the posterior $p({\theta_{j}},G\mid{z_{j}})$ of the DP mixture. Posterior estimation is conducted via the memory-based Variational Bayes (memoVB) [hughes2013memoized] algorithm, which supports incremental learning by sharing global statistics across tasks.
During VLA training, our task-centric representation learning and knowledge space updating are jointly performed in a fully self-supervised manner. The perception encoder continuously provides latent features to update the knowledge distribution, while the knowledge space, via KL loss $L_{kl}$, constrains $z_{j}$ to infer under task or skill hypotheses. This design establishes a self-evolving cycle that autonomously retains and discovers knowledge over time.
Table 1: Overall performance on the LIBERO benchmark. Baselines are grouped into pretrained and scratch categories. Results are reported in percentages. The best scratch model is shown in bold, and the best overall in underline.
tasks metrics UniVLA [UniVLA] MoDE [MoDE] MoDE [MoDE] T-Stellar TS-Stellar
(fine-tuned) (fine-tuned) (scratch) (scratch, ours) (scratch, ours)
LIBERO-goal FWT (↑) 87.4 69.7 68.5 81.9 78.5
NBT (↓) 59.6 27.9 33.6 19.3 21.9
AUC (↑) 39.9 43.3 38.5 63.2 58.3
Final SR (↑) 25.2 37.5 36.0 66.8 64.3
LIBERO-long FWT (↑) 70.6 68.3 66.6 78.6 74.7
NBT (↓) 32.38 45.7 45.4 44.3 37.9
AUC (↑) 44.97 30.6 28.4 41.2 41.7
Final SR (↑) 17.0 18.7 18.6 31.9 36.8
LIBERO-30 * FWT (↑) 87.9 52.9 48.2 79.6 73.3
NBT (↓) 57.7 11.9 15.4 31.0 27.8
AUC (↑) 36.5 43.0 33.6 49.7 48.3
Final SR (↑) 29.6 40.3 28.5 42.9 42.6
Table 2: Performance on real-world experiments. MoDE is our modified MoDE [MoDE] for dual arm tasks, and all models are trained from scratch. The best performance is highlighted in bold. All models are trained from scratch.
tasks metrics MoDE [MoDE] T-Stellar TS-Stellar
Dual arm real world task FWT (↑) 96.7 96.7 96.7
NBT (↓) 45.0 35.0 21.7
AUC (↑) 69.4 76.1 82.2
Final SR (↑) 46.7 66.7 83.3
3.4 Action-Knowledge-guided Expert Routing
Figure 2: Knowledge-prior-routed MoE action head. Two knowledge embeddings, relation and top-K semantic, are computed for expert routing, alongside language, noise, observation and noise action tokens fed into the denoising transformer.
After learning the knowledge space and task-centric latents, we introduce a knowledge-routed MoE module to allocate task-specific action prediction parameters, minimizing interference between unrelated tasks while enhancing sharing among related ones. A diffusion-based MoE action head is adopted referring to MoDE [MoDE], replacing linear layers in the denoising transformer with MoE modules:
$\text{MoE}(tok,\mathbf{f}_{ro})=\sum_{i=1}^{N_{e}}\mathbf{Router}_{i}(\mathbf{f}_{ro})\cdot\mathbf{Expert}_{i}(tok)$ (7)
Here, toktok is the input token feature tensor and $\mathbf{f}_{ro}$ the routing features, defined in MoDE as the noise embedding $\mathbf{f}_{router}={tok}_{noise}=\phi(\sigma_{\mathbf{t}})$. Unlike MoDE’s noise-level routing, we guide expert allocation via the knowledge space, enabling continual task learning with task-specific parameter sharing and differentiation. Since DPMM and HDP produce variable numbers of clusters, directly feeding their task distributions into a fixed-dimension network is impractical. Thus, we design two knowledge embedding formats, as shown in Fig. 2.
Knowledge Relation Embedding encodes the distance between latent $z$ and cluster centers $\mu_{k}$. Given the membership probability $p_{k}$ of $z$ for cluster $k$ as described in Sec. 3.3, relation embedding is computed as:
$\mathbf{f}_{\text{R}}=\sum_{k=1}^{K_{c}}p_{k}\,|z-\mu_{k}|$ (8)
Top-K Semantic Embedding represents the discrete skill semantic affiliation of $z$ by aggregating learnable embeddings of the top-K assigned clusters:
$\mathbf{f}_{\text{S}}=\sum_{k\in\mathcal{I}_{T}}p_{k}\cdot\text{Embed}(k),\mathcal{I}_{T}=\text{TopKIndice}({p_{1},\dots,p_{K}},z)$ (9)
Finally, the embedded knowledge of current task and skill could be represented as:
$\mathbf{f}_{\text{know}}=\left[\,z\,\|\mathbf{f}_{\text{R}}\;\|\;\mathbf{f}_{\text{S}}\,\right]$ (10)
The knowledge embedding $\mathbf{f}_{\text{know}}$ is introduced as an additional guidance token ${tok}_{know}$, which, together with noise token ${tok}_{noise}$ and language token ${tok}_{lang}$, are concatenated to feed the router for task-aware expert allocation.
By incorporating knowledge space guidance into expert routing, the model could achieve effective CIL through balanced parameter sharing and isolation, enabling related tasks to benefit from shared experience while preventing catastrophic forgetting across dissimilar ones.
4 Experiments
Our experiments address four questions: (I) Can T-Stellar and TS-Stellar surpass state-of-the-art VLA models in end-to-end continual learning? (II) Does TS-Stellar’s hierarchical task–skill knowledge space outperform T-Stellar’s task-only modeling? (III) Do DP-based knowledge space and knowledge-guided expert routing enhance continual learning? (IV) How does the DP-based knowledge space gradually discover and preserve skills?
4.1 Simulation Experiments
We evaluate our models on the LIBERO benchmark [liu2023libero], including LIBERO-goal, LIBERO-long, and LIBERO-30*, a 30-task subset of LIBERO-90 constructed by ourselves, for comprehensive evaluation under diverse goals, long-horizon reasoning, and a large number of tasks. Demo data consist of static and wrist camera images, and actions include end-effector poses and gripper states.
For continual learning, models are trained sequentially following the task order in LIBERO, using Experience Replay with only 1% of past data stored, which suffices for LIBERO tasks while reducing memory overhead. Both T-Stellar and TS-Stellar are trained from scratch.
Baselines. Since no VLA has been evaluated under CIL, we select strong multitask performers on LIBERO and re-train them under this setting for fair comparison. The baselines include: 1) Pretrained UniVLA [UniVLA], which learns task-centric representations via large-scale latent action pretraining and achieves state-of-the-art multitask results; 2) MoDE [MoDE], a diffusion-based Mixture-of-Experts VLA similar to our action head, serving as a key comparison. Both pretrained and scratch variants are evaluated, consistent with its original setups.
Metrics. We evaluate continual learning using four metrics: FWT (forward transfer), NBT (negative backward transfer), AUC (area under the success rate curve), [IsCiL, wan2024lotus, diaz2018don, liu2023libero] and Final SR (final average success rate), all computed from the success rate matrix $R$ where $R_{i,j}$ is the success on task $j$ after learning $i$ tasks. FWT measures forward learning, NBT measures forgetting, AUC reflects overall performance stability, and Final SR is overall success after training all tasks. Formally, $\text{FWT}=\frac{1}{M}\sum_{m=1}^{M}R_{m,m}$, $\text{NBT}=\frac{1}{M}\sum_{m=1}^{M}\frac{1}{M-m}\sum_{q=m+1}^{M}(R_{m,m}-R_{q,m})$, $\text{AUC}=\frac{1}{M}\sum_{m=1}^{M}\frac{1}{M-m+1}\left(R_{m,m}+\sum_{q=m+1}^{M}R_{q,m}\right)$, $\text{Final SR}=\frac{1}{M}\sum_{j=1}^{M}R_{M,j}$. Each policy is evaluated on all previous tasks 100 times to form the SR matrix.
Results. Tab. 1 shows that T-Stellar and TS-Stellar achieve the highest final success rates across all scenarios, with over 50% average improvement relative to both pretrained and from-scratch baselines [UniVLA, MoDE]. They also outperform most baselines in AUC and NBT, demonstrating strong and stable continual learning performance.
Among baselines, UniVLA is the most competitive, benefiting from its large model capacity and unified action space, which yield strong FWT and long-horizon performance. However, without explicit task-specific knowledge and task-guided action allocation, it suffers from poor retention in single-scene and many-task scenarios, often showing highest NBT and lowest Final SR, highlighting the instability of current VLA models under continual learning.
Figure 3: T-SNE visualization of Stellar VLA Latent representations after 1, 4, 6, 8, and 10 tasks on LIBERO-long are shown. Task names are abbreviated for clarity. T-Stellar models discrete task distributions, and TS-Stellar learn relevant skill across tasks.
Table 3: Knowledge space ablation experiments on LIBERO and real-world tasks. The best performance is shown in bold.
tasks metrics w/o VAE w/o KS T-Stellar TS-Stellar
LIBERO-goal FWT (↑) 70.8 76.6 81.9 78.5
NBT (↓) 30.1 21.9 19.3 21.9
AUC (↑) 43.5 56.4 63.2 58.3
Final SR (↑) 36.2 43.8 66.8 64.3
LIBERO-long FWT (↑) 60.9 73.5 78.6 74.7
NBT (↓) 38.9 47.0 44.3 37.9
AUC (↑) 27.2 34.5 41.2 41.7
Final SR (↑) 17.1 22.0 31.9 36.8
LIBERO-30 * FWT (↑) 45.0 45.1 79.6 73.3
NBT (↓) 12.0 4.9 31.0 27.8
AUC (↑) 32.7 38.9 49.7 48.3
Final SR (↑) 26.3 35.8 42.9 42.6
Dual arm real world task FWT (↑) 90.0 96.7 96.7 96.7
NBT (↓) 55.0 58.3 31.0 27.8
AUC (↑) 58.9 63.3 76.1 82.2
Final SR (↑) 36.7 40.0 66.7 83.3
Compared with both fine-tuned and from-scratch MoDE, our methods achieve the best FWT and AUC across all settings. MoDE variants achieve better NBT on LIBERO-30, mainly because their FWT is only about half of ours, resulting in lower retention pressure. This reflects the inherent trade-off between strong forward transfer and maintaining low backward transfer in continual learning.
Therefore, in response to Question (I), although trained from scratch, our methods consistently outperform both pretrained and from-scratch VLA models in overall continual learning. Notably, TS-Stellar surpasses T-Stellar on long-horizon tasks but slightly lags on simpler tasks, likely because the advantages of hierarchical task–skill modeling are less pronounced in such settings.
4.2 Real-Word Experiment
To further evaluate real-world effectiveness, we conduct dual-arm manipulation experiments on a humanoid robot platform. Three tasks are designed as a continual learning sequence: 1) “Transfer Magic Stick”, 2) “Pick up Bag”, and 3) “Handover Toy”. Models are trained sequentially with a 5% replay rate for stability. Observations are collected from a head camera and dual wrist cameras, while outputs consist of joint poses and gripper states. Evaluation follows Sec. 4.1, with each task tested ten times.
Baselines. As mentioned, MoDE [MoDE] serves as a key reference for our method. However, since MoDE was originally designed for single-arm Cartesian action prediction, we modified its architecture without altering core components to ensure fair comparison on dual-arm joint-space tasks. The modified trained from scratch version is denoted as MoDE, and its pretrained model is thus not used.
Results.
Figure 4: Behavior visualization on “Pick up Bag” after training on “Handover Toy”. TS-Stellar achieves the most synchronized dual-arm motion; T-Stellar hesitates slightly; MoDE and w/o KS show strong desynchronization and ultimately fail.
FWT, NBT, AUC and Final SR are reported in Tab. 2. The results show that T-Stellar and TS-Stellar substantially outperform MoDE in NBT, AUC, and Final SR while achieve strong FWT, demonstrating superior continual learning on dual-arm real-world tasks. TS-Stellar further surpasses T-Stellar across all these metrics, suggesting that task–skill modeling is particularly effective for long-horizon skill sequences and complex action mappings, thus addressing Question (II).
Table 4: Prior-guided expert routing ablation experiments on LIBERO-long. The best performance is highlighted in bold.
tasks metrics T-Stellar T-Stellar TS-Stellar TS-Stellar
w/o K-route w/o K-route
LIBERO-long FWT (↑) 74.4 78.6 76.1 74.7
NBT (↓) 44.1 44.3 42.7 37.9
AUC (↑) 38.0 41.2 39.4 41.7
Final SR (↑) 20.9 31.9 30.8 36.8
4.3 Ablation Studies
In this section, we conduct two sets of ablation studies to investigate the effectiveness of the proposed knowledge space and expert routing mechanisms.
Knowledge Space Ablation. We perform ablations by replacing the DP-based knowledge space with a standard Gaussian $(\mu=0,\sigma=1)$ (w/o KS) and by removing latent task-centric representation learning (w/o VAE). Experiments are conducted across all simulation and real-world tasks, as shown in Tab. 3. Results demonstrate that our DP-based knowledge spaces significantly improve all four metrics—except for NBT on LIBERO-30, which is affected by weaker forward learning of the ablations. Overall, the results validates the effectiveness of our DPMM-based task-centric and HDP-based task–skill knowledge spaces.
Knowledge Guidance for Expert Routing Effectiveness. We evaluate T-Stellar and TS-Stellar on LIBERO-long without the knowledge-guidance token ${tok}_{know}$, as shown in Tab. 4. Both models achieve higher AUC and Final SR with knowledge guidance.
Through ablation studies, we could answer Question (III): The DP-based knowledge space enhances continual learning by capturing task and skill distributions for better task separation and knowledge sharing, while knowledge-routed MoE further improves stability and performance through dynamic expert allocation.
4.4 Visualization for Continual Learning process
To demonstrate how the proposed method support continual skill evolution, we visualize the learning process.
Knowledge Evolution During Continual Learning. We analyze knowledge space evolution on LIBERO-long. Fig. 3 shows t-SNE plots of latent representations after training on selected tasks. T-Stellar learns distinct clusters for different tasks, and TS-Stellar connects tasks through shared subskills. For example, deep green and light green tasks overlaps since they share the same “put moka on stove” subgoal.
Execution Behavior Analysis. To understand performance drops during continual learning, we visualize real-world robot actions. Fig. 4 shows the “Pick up Bag” task after “Handover Toy” training for T-Stellar, TS-Stellar, MoDE, and w/o KS ablation. For the “Pick up Bag” task, both grippers close synchronously, but asynchrony could emerge after training on “Handover Toy”. TS-Stellar maintains stable, synchronized gripper and arm motion. T-Stellar shows mild hesitation yet still lifts the bag. In contrast, MoDE and w/o KS exhibit clear gripper desynchronization; MoDE fails from over-squeezing the bag, while w/o KS fails by rigidly following the action pattern of task “Handover Toy” .
Overall, visualization and behavior analysis answer Question (IV): T-Stellar learns distinct clusters for new tasks, whereas TS-Stellar captures inter-task skill relations, forming an overlapping task–skill knowledge space. Retaining skill priors enables stable old-task execution and mitigates behavioral drift during continual learning.
5 Conclusion
We introduce Stellar VLA, a knowledge space driven policy for continual imitation learning, with two variants: T-Stellar, which constructs task-centric knowledge spaces, and TS-Stellar, which models hierarchical task–skill relations. By leveraging the evolving knowledge space to guide expert selection, our models reduce interference between tasks while enabling continual skill acquisition. Experiments on LIBERO and real-world tasks demonstrate significant performance gains, and visualizations show the effective evolution of task and skill knowledge. Future work could explore larger skill spaces to scale up continual learning for larger VLA models.